what is Chapman model

Grade:12

1 Answers

Gman Namg
8 years ago
The Chapman model accounts for ozone formation and destruction in an atmosphere that contains only O2
Initiation: O2 + hν →O + O 185 nm
Propagation: O + O2 +M→O3 + M* ΔrH = −106.6 kJ mol−1 v = k2[O][O2][M]
O3 + hν →O2 + O 210 nm
Termination: O + O3→O2 + O2 ΔrH = −391.9 kJ mol−1 v = k4[O][O3]
O + O +M→O2 + M* v = k5[O]2[M]
where M is an arbitrary third body, such as O2 in an ‘oxygen-only’ atmosphere, which
helps to remove excess energy from the products of combination and recombination
reactions. The mechanism shows that absorption of radiation by O2 and O3 during
the daytime leads to the production of reactive O atoms, which, in turn, participate in
exothermic reactions that are responsible for the heating of the stratosphere.
approve if useful